# Preference Elicitation for Offline Reinforcement Learning

### v1

@@ 5 Preference Elicitation from Offline Trajectories
OPRL Uniform Sampling: $\tau_{1},\tau_{2}\sim\mathcal{D}_{offline}$
OPRL Uncertainty Sampling: $\tau_{1},\tau_{2}=\text{argmax}_{\tau_{1},\tau_{2}\in\mathcal{D}_{offline}}u_{P_{R}}(\tau_{1},\tau_{2})$
@@
We obtain the following result, demonstrated in Section A.4. The suboptimality of the estimated policy $\hat{\pi}^{*}$ is bounded by the policy evaluation error for the optimal policy $\pi^{*}$. This error decomposes into a term depending on transition model estimation, and one on reward model estimation.
@@
For any $\delta\in(0,1]$, let $\beta_{T}=c_{T}^{\prime}\log(H\mathcal{N}_{\mathcal{F}_{T}}(1/N_{o})/\delta)/N_{o}$ and $\beta_{R}={c_{R}^{\prime}\log(\mathcal{N}_{\mathcal{F}_{R}}(1/N_{p})/\delta)}/{N_{p}}$, where $N_{o}=H|\mathcal{D}_{offline}|$ is the number of observed transitions in the observational dataset and $c_{T}^{\prime},c_{R}^{\prime}$ are universal constants. The policy $\hat{\pi}^{*}$ estimated by Algorithm 1, with preference elicitation based on offline trajectories, achieves the following suboptimality with probability $1-\delta$:
$V^{\pi^{*}}-V^{\hat{\pi}^{*}}\leq\underbrace{HR_{max}C_{T}(\mathcal{F}_{T},\pi^{*})\sqrt{\frac{c_{T}\log(H\mathcal{N}_{\mathcal{F}_{T}})(1/N_{o})/\delta)}{N_{o}}}}_{\text{transition term }\epsilon_{T}}+\underbrace{2\alpha\kappa C_{R}(\mathcal{F}_{R},\pi^{*})\sqrt{\frac{c_{R}\log(\mathcal{N}_{\mathcal{F}_{R}}(1/N_{p})/\delta)}{N_{p}}}}_{\text{reward term}},$
where $\alpha=1$ for uniform sampling or $\alpha\leq 1$ for uncertainty sampling, $C_{R}$ is a concentrability measure for the reward function, $\kappa=\sup_{r\in[-R_{max},R_{max}]}\frac{1}{\sigma^{\prime}(r)}$ measures the degree of non-linearity of the sigmoid function, and $c_{T},c_{R}$ are universal constants.
In the special case where both the transition and reward functions are learned on a fixed initial preference dataset (no preference elicitation; $|\mathcal{D}_{offline}|=2N_{p}$), we recover Theorem 1 from Zhan et al. [2023a]. Importantly, parameter $\alpha$ allows us to motivate the superior efficiency of uncertainty sampling over uniform sampling, observed empirically in Shin et al. [2022] and in our own experiments (Section 7). Uncertainty sampling learns accurate reward models with fewer preference queries when $\alpha<1$, but can perform like uniform sampling in harder problems ($\alpha=1$).
@@ 6 Preference Elicitation from Simulated Trajectories
We now propose our alternative strategy for generating trajectories for offline preference elicitation: Simulated Offline Preference-based Reward Learning (Sim-OPRL). This method simulates trajectories $(\tau_{1},\tau_{2})$ by leveraging the learned environment model. This overcomes a limitation of OPRL, which is only designed to reduce uncertainty about the reward functions in $\mathcal{R}$, by instead reducing uncertainty about which policies are plausibly optimal. Our approach is inspired by efficient online preference elicitation algorithms [Saha et al., 2023, Chen et al., 2022], which we modify for practical implementation. We account for the offline nature of our problem by avoiding regions out of the distribution of the data: the sampling strategy is optimistic with respect to uncertainty in rewards, but pessimistic with respect to uncertainty in transitions.
We summarize our approach to generating simulated trajectories for preference elicitation in Algorithm 2 and refer the reader to Appendix B for practical implementation details. First, we construct a set of candidate optimal policies $\Pi_{offline}$, containing policy $\pi^{*}_{offline}$ (optimal under the pessimistic model and the true reward function) with high probability – as demonstrated in Section A.5.2. Next, within this set of candidate policies, we identify the two most exploratory policies $\pi_{1},\pi_{2}$, chosen to maximize preference uncertainty $u_{P_{R}}$. Finally, we roll out these policies within our learned transition model to generate a trajectory pair $(\tau_{1},\tau_{2})$ for preference feedback.
@@
1:Pessimistic transition model $\hat{T}_{inf}$. Reward confidence set $\mathcal{R}$ and preference uncertainty function $u_{P_{R}}$.
@@
3:Estimate optimal offline policy set: $\Pi_{offline}=\{\pi\>|\>\pi=\text{argmax}_{\pi\in\Pi}\mathbb{E}_{\tau\sim d^{\pi}_{\hat{T}_{inf}}}\big{[}\tilde{R}(\tau)\big{]}\>\forall\tilde{R}\in\mathcal{R}\}$.
4:Identify exploratory policies: $\pi_{1},\pi_{2}=\text{argmax}_{\pi_{1},\pi_{2}\in\Pi_{offline}}\mathbb{E}_{\tau_{1}\sim d^{\pi_{1}}_{\hat{T}_{inf}},\tau_{2}\sim d^{\pi_{2}}_{\hat{T}_{inf}}}\left[u_{P_{R}}(\tau_{1},\tau_{2})\right]$
5:Rollouts in model: $\tau_{1}\sim d^{\pi_{1}}_{\hat{T}_{inf}},\tau_{2}\sim d^{\pi_{2}}_{\hat{T}_{inf}}$.
@@ 6.1 Theoretical Guarantees
We decompose suboptimality in a similar way to Section 5.1, but obtain a reward suboptimality term that depends on the learned dynamics model instead of the true one, and on $\pi^{*}_{offline}$ instead of $\pi^{*}$:
$V^{\pi^{*}}-V^{\hat{\pi}^{*}}\leq\underbrace{(V^{{\pi}^{*}}_{T,R}-V^{\pi^{*}}_{\hat{T}_{inf},R})}_{\text{transition term }\epsilon_{T}}+\underbrace{(V^{\pi^{*}_{offline}}_{\hat{T}_{inf},R}-V^{\pi^{*}_{offline}}_{\hat{T}_{inf},\hat{R}_{inf}})}_{\text{reward term}}.$ (3)
Analysis of the suboptimality due to transition error is identical to above, but the reward term is thus significantly different. By design, our sampling strategy ensures good coverage of preferences over $\pi^{*}_{offline}$ within the learned environment model, which eliminates the concentrability term for the reward $C_{R}$. We refer the reader to Section A.5 for the proof of Theorem 6.1.
@@
For any $\delta\in(0,1]$, let $\beta_{T}=c_{T}^{\prime}\log(H\mathcal{N}_{\mathcal{F}_{T}}(1/N_{o})/\delta)/N_{o}$ and $\beta_{R}={c_{R}^{\prime}\log(\mathcal{N}_{\mathcal{F}_{R}}(1/N_{p})/\delta)}/{N_{p}}$, where $N_{o}=H|\mathcal{D}_{offline}|$ is the number of observed transitions in the observational dataset and $c_{T}^{\prime},c_{R}^{\prime}$ are universal constants. The policy $\hat{\pi}^{*}$ estimated by Algorithm 1, with a preference sampling strategy based on rollouts in the learned transition model, achieves the following suboptimality with probability $1-\delta$:
$V^{\pi^{*}}-V^{\hat{\pi}^{*}}\leq HR_{max}C_{T}(\mathcal{F}_{T},\pi^{*})\sqrt{\frac{c_{T}\log(H\mathcal{N}_{\mathcal{F}_{T}})(1/N_{o})/\delta)}{N_{o}}}+2\kappa\sqrt{\frac{c_{R}\log(\mathcal{N}_{\mathcal{F}_{R}}(1/N_{p})/\delta)}{N_{p}}}.$
where $\kappa=\sup_{r\in[-R_{max},R_{max}]}\frac{1}{\sigma^{\prime}(r)}$ measures the degree of non-linearity of the sigmoid function, and $c_{T},c_{R}$ are universal constants.
@@ 6.2 Discussion
Our theoretical results demonstrate that the learned policy can achieve performance comparable to the optimal policy, and thus satisfy our optimality criterion in Definition 3.1, provided it is covered by the offline data ($C_{T}(\mathcal{F}_{T},\pi^{*}),C_{R}(\mathcal{F}_{R},\pi^{*})<\infty$). Empirical results in Section 7 confirm that performance is poor when the behavioral policy is suboptimal, inducing a large $C_{T}$ or $C_{R}$.
@@
Offline Trajectories vs. Simulated Rollouts. While both OPRL and Sim-OPRL depend on the offline dataset for estimating environment dynamics, they induce different suboptimality in modeling preference feedback. Simulated rollouts are designed to achieve good coverage of the optimal offline policy ${\pi}^{*}_{offline}$, which avoids wasting preference budget on trajectories with low rewards or high transition uncertainty. In contrast, as shown in Zhan et al. [2023a], due to the dependence of preferences on full trajectories, the reward concentrability term $C_{R}$ in Theorem 5.1 can be very large. An advantage of sampling from the offline buffer, however, is that it is not sensitive to the quality of the model.
@@
Transition vs. Preference Model Quality. Our theoretical analysis also suggests an interesting trade-off in the sample efficiency of our approach, depending on the accuracy of the transition model. The width of the confidence interval reduces as significance parameter $\delta$ or dataset size increase, or as function class complexity $\mathcal{N}_{\mathcal{F}_{T}}$ decreases. For a target suboptimality gap $\epsilon$, provided the optimal offline policy $\pi^{*}_{offline}$ has a gap $\epsilon_{T}<\epsilon$, then the number of preferences required is of the order of $\mathcal{O}({\log(1/\delta)}/{(\epsilon-\epsilon_{T})^{2}})$. A more accurate transition model should therefore require fewer preference samples to achieve a given suboptimality, which we again confirm empirically.
@@ 7 Experimental results
Figure 1: Star MDP.
We demonstrate the effectiveness of preference elicitation for offline reinforcement learning in practice and compare the different sampling strategies introduced in Sections 5 and 6: OPRL with uniform and uncertainty-sampling, and Sim-OPRL. Our experiments explore both simple MDPs and more complex decision-making environments.
@@
Baselines. For comparison, we also propose a practical implementation of Preference-based Optimistic Planning (PbOP), an uncertainty-based preference elicitation approach over trajectory rollouts in the true environment [Chen et al., 2022]. Finally, we report the performance of $\pi^{*}_{offline}$ and $\pi^{*}$ as upper bounds for the performance of our algorithm: the former is trained in the learned transition model with access to the true reward, and the latter has full knowledge of both transition and reward function. We refer the reader to Appendix B for implementation details.
Star MDP. First, consider the simple tabular MDP illustrated in Figure 1, consisting of four states branching from a central state. We defer transition and reward details to Appendix C. Preferences collected over samples from the offline dataset learn slowly about the negative reward in the bottom state, as it is always included in the sampled trajectories. Instead, simulated rollouts can query a direct comparison between the optimal path and one that includes it. We thus find in Figure 2 that our preference elicitation strategy based on simulated rollouts achieves better environment returns than OPRL approaches which sample from the offline buffer, with much fewer preference queries.
This example also illustrates the importance of pessimism with respect to the transition model. Even with access to true rewards, $\pi^{*}_{offline}$ is pessimistic to avoid the out-of-distribution state, as it is unclear how to reach it. Thus, in Figure 2(b), we see a drop in performance if pessimism is not applied to the output policy (purple lines). This confirms the theoretical insights from Zhu et al. [2023], Zhan et al. [2023a], who demonstrate the importance of pessimism in offline preference-based RL problems. Pessimism is also crucial in simulated rollouts, to avoid wasting preference budget on regions of low confidence — as value estimates are inaccurate in any case. This is reflected in the lower sample efficiency of model rollouts which do not apply pessimism with respect to the transition model in Figure 2(b) (brown line), and which could be seen as the naive adaptation of online preference elicitation methods to our setting [Chen et al., 2022, Lindner et al., 2022]. We also note the importance of optimism with respect to the reward uncertainty, both in OPRL in Figure 2(a) and in our own model-based rollouts in Figure 2(b).
@@
Finally, as an upper bound for the performance of our algorithm, we include baselines that have access to the environment in Figure 2(a): we report the performance of the optimal policy $\pi^{*}$, as well as that of an algorithm querying feedback over optimistic rollouts in the real environment [Chen et al., 2022, PbOP]. Final environment returns are higher than with our method, as they do not suffer from the limited coverage of the transition model. As supported by our theoretical analysis, this result stresses the importance of having a high-quality transition model to make our method effective. We explore this in more detail in the following experiment.
(a) Comparison against baselines.
(b) Algorithm Ablations.
@@
Figure 2: Empirical results on the Star MDP. Mean and 95% confidence interval over 20 experiments. Environment returns are normalized between 0 and 100. Only OPRL and Sim-OPRL are fully offline, all other methods have access to either environment interaction and/or to the true reward function.
@@
Transition vs. Preference Model Quality. Next, we empirically study the trade-off between transition and preference model performance in our problem setting. In the low-data regime, the error $\epsilon_{T}$ incurred in estimating the value function due to the misspecification of the transition model is large. As dictated by our theoretical analysis and as visualized in Figure 3(a), this significantly increases the number of preference samples $N_{p}$ required to achieve good final performance. At the other end of the spectrum, if the offline dataset is large and allows modeling the transition model accurately, then $\epsilon_{T}$ is small and the number of preference samples $N_{p}$ needed shrinks. We observe a similar trend for both Sim-OPRL and our OPRL uncertainty-sampling baseline.

### v2

@@ 5 Preference Elicitation from Offline Trajectories
OPRL Uniform Sampling: $\tau_{1},\tau_{2}\sim\mathcal{D}_{\textrm{offline}}$
OPRL Uncertainty Sampling: $\tau_{1},\tau_{2}=\text{argmax}_{\tau_{1},\tau_{2}\in\mathcal{D}_{\textrm{offline}}}u_{P_{R}}(\tau_{1},\tau_{2})$
@@
We obtain the following result, proven in Section A.4. The suboptimality of the estimated policy $\hat{\pi}^{*}$ is bounded by the policy evaluation error for the optimal policy $\pi^{*}$. This error decomposes into a term depending on transition model estimation, and one on reward model estimation.
@@
For any $\delta\in(0,1]$, let $\beta_{T}=c^{\textrm{MLE}}_{T}\log(H\mathcal{N}_{\mathcal{F}_{T}}(1/N_{o})/\delta)/N_{o}$ and $\beta_{R}={c^{\textrm{MLE}}_{R}\log(\mathcal{N}_{\mathcal{F}_{R}}(1/N_{p})/\delta)}/{N_{p}}$, where $N_{o}=H|\mathcal{D}_{\textrm{offline}}|$ is the number of observed transitions in the observational dataset and $c^{\textrm{MLE}}_{T},c^{\textrm{MLE}}_{R}$ are universal constants. The policy $\hat{\pi}^{*}$ estimated by Algorithm 1, with preference elicitation based on offline trajectories, achieves the following suboptimality with probability $1-\delta$:
$V^{\pi^{*}}-V^{\hat{\pi}^{*}}\leq\underbrace{HR_{\textrm{max}}C_{T}(\mathcal{F}_{T},\pi^{*})\sqrt{\frac{c_{T}}{N_{o}}\log\left(\frac{H}{\delta}\mathcal{N}_{\mathcal{F}_{T}}\left(\frac{1}{N_{o}}\right)\right)}}_{\text{transition term }\epsilon_{T}}+\underbrace{2\alpha\kappa C_{R}(\mathcal{F}_{R},\pi^{*})\sqrt{\frac{c_{R}}{N_{p}}\log\left(\frac{1}{\delta}\mathcal{N}_{\mathcal{F}_{R}}\left(\frac{1}{N_{p}}\right)\right)}}_{\text{reward term}},$
where $\alpha=1$ for uniform sampling or $\alpha\leq 1$ for uncertainty sampling, $C_{R}$ is a concentrability measure for the reward function, $\kappa=\sup_{r\in[-R_{\textrm{max}},R_{\textrm{max}}]}\frac{1}{\sigma^{\prime}(r)}$ measures the degree of non-linearity of the link function, and $c_{T},c_{R}$ are universal constants.
In the special case where both the transition and reward functions are learned on a fixed initial preference dataset (no preference elicitation; $|\mathcal{D}_{\textrm{offline}}|=2N_{p}$), we recover Theorem 1 from Zhan et al. [2023a]. Importantly, the coefficient $\alpha$ allows us to motivate the superior efficiency of uncertainty sampling over uniform sampling, observed empirically in Shin et al. [2022] and in our own experiments (Section 7). Uncertainty sampling learns accurate reward models with fewer preference queries when $\alpha<1$, but can perform like uniform sampling in harder problems ($\alpha=1$).
@@ 6 Preference Elicitation from Simulated Trajectories
We now propose our alternative strategy for generating trajectories for offline preference elicitation: Simulated Offline Preference-based Reward Learning (Sim-OPRL). This method simulates trajectories $(\tau_{1},\tau_{2})$ by leveraging the learned environment model. This overcomes a limitation of OPRL, which is only designed to reduce uncertainty about the reward functions in $\mathcal{R}$, by instead reducing uncertainty about which policies are plausibly optimal. Our approach is inspired by efficient online preference elicitation algorithms [Saha et al., 2023, Chen et al., 2022], which we modify for practical implementation. We account for the offline nature of our problem by avoiding regions that are out of the distribution of the data: the sampling strategy is optimistic with respect to uncertainty in rewards, but pessimistic with respect to uncertainty in transitions.
We summarize our approach to generating simulated trajectories for preference elicitation in Algorithm 2. First, we construct a set of candidate optimal policies $\Pi_{\textrm{offline}}$, containing policy $\pi^{*}_{\textrm{offline}}$ (optimal policy under the pessimistic model and the true reward function) with high probability – as demonstrated in Section A.5.2. Next, within this set of candidate policies, we identify the two most exploratory policies $\pi_{1},\pi_{2}$, chosen to maximize preference uncertainty $u_{P_{R}}$. Finally, we roll out these policies within our learned transition model to generate a trajectory pair $(\tau_{1},\tau_{2})$ for preference feedback.
@@
1:Pessimistic transition model $\hat{T}_{\textrm{inf}}$. Reward confidence set $\mathcal{R}$ and preference uncertainty function $u_{P_{R}}$.
@@
3:Estimate optimal offline policy set: $\Pi_{\textrm{offline}}=\{\pi\>|\>\exists\tilde{R}\in\mathcal{R}:\pi=\text{argmax}_{\pi\in\Pi}\mathbb{E}_{\tau\sim d^{\pi}_{\hat{T}_{\textrm{inf}}}}\big{[}\tilde{R}(\tau)\big{]}\}$
4:Identify exploratory policies: $\pi_{1},\pi_{2}=\text{argmax}_{\pi_{1},\pi_{2}\in\Pi_{\textrm{offline}}}\mathbb{E}_{\tau_{1}\sim d^{\pi_{1}}_{\hat{T}_{\textrm{inf}}},\tau_{2}\sim d^{\pi_{2}}_{\hat{T}_{\textrm{inf}}}}\left[u_{P_{R}}(\tau_{1},\tau_{2})\right]$
5:Rollouts in model: $\tau_{1}\sim d^{\pi_{1}}_{\hat{T}_{\textrm{inf}}},\tau_{2}\sim d^{\pi_{2}}_{\hat{T}_{\textrm{inf}}}$.
We first provide a theoretical analysis of the performance of Sim-OPRL, before proposing a practical implementation of our entire preference elicitation and policy optimization algorithm.
@@ 6.1 Theoretical Guarantees
We decompose suboptimality in a similar way to Section 5.1, but obtain a reward suboptimality term that depends on the learned dynamics model instead of the true one, and on $\pi^{*}_{\textrm{offline}}$ instead of $\pi^{*}$:
$V^{\pi^{*}}-V^{\hat{\pi}^{*}}\leq\underbrace{(V^{{\pi}^{*}}_{T,R}-V^{\pi^{*}}_{\hat{T}_{\textrm{inf}},R})}_{\text{transition term }\epsilon_{T}}+\underbrace{(V^{\pi^{*}_{\textrm{offline}}}_{\hat{T}_{\textrm{inf}},R}-V^{\pi^{*}_{\textrm{offline}}}_{\hat{T}_{\textrm{inf}},\hat{R}_{\textrm{inf}}})}_{\text{reward term}}.$ (4)
Analysis of the suboptimality due to transition error is identical to above, but the reward term is thus significantly different. By design, our sampling strategy ensures good coverage of preferences over $\pi^{*}_{\textrm{offline}}$ within the learned environment model, which eliminates the concentrability term for the reward $C_{R}$. We refer the reader to Section A.5 for the proof of Theorem 6.1.
@@
For any $\delta\in(0,1]$, let $\beta_{T}=c^{\textrm{MLE}}_{T}\log(H\mathcal{N}_{\mathcal{F}_{T}}(1/N_{o})/\delta)/N_{o}$ and $\beta_{R}={c^{\textrm{MLE}}_{R}\log(\mathcal{N}_{\mathcal{F}_{R}}(1/N_{p})/\delta)}/{N_{p}}$, where $N_{o}=H|\mathcal{D}_{\textrm{offline}}|$ is the number of observed transitions in the observational dataset and $c^{\textrm{MLE}}_{T},c^{\textrm{MLE}}_{R}$ are universal constants. The policy $\hat{\pi}^{*}$ estimated by Algorithm 1, with a preference sampling strategy based on rollouts in the learned transition model, achieves the following suboptimality with probability $1-\delta$:
$V^{\pi^{*}}-V^{\hat{\pi}^{*}}\leq\underbrace{HR_{\textrm{max}}C_{T}(\mathcal{F}_{T},\pi^{*})\sqrt{\frac{c_{T}}{N_{o}}\log\left(\frac{H}{\delta}\mathcal{N}_{\mathcal{F}_{T}}\left(\frac{1}{N_{o}}\right)\right)}}_{\text{transition term }\epsilon_{T}}+\underbrace{2\kappa\sqrt{\frac{c_{R}}{N_{p}}\log\left(\frac{1}{\delta}\mathcal{N}_{\mathcal{F}_{R}}\left(\frac{1}{N_{p}}\right)\right)}}_{\text{reward term}}.$
where $\kappa=\sup_{r\in[-R_{\textrm{max}},R_{\textrm{max}}]}\frac{1}{\sigma^{\prime}(r)}$ measures the degree of non-linearity of the link function, and $c_{T},c_{R}$ are universal constants.
@@ 6.2 Discussion
Our theoretical results demonstrate that the learned policy can achieve performance comparable to the optimal policy, and thus satisfy our optimality criterion in Definition 3.1, provided it is covered by the offline data ($C_{T}(\mathcal{F}_{T},\pi^{*}),C_{R}(\mathcal{F}_{R},\pi^{*})<\infty$). Following our analysis, a suboptimal dataset requires more preferences to achieve a certain policy performance, as the concentrability terms $C_{T}$ or $C_{R}$ are large. Empirical results in Section 7 confirm that sample efficiency is worse when the behavioral policy is more suboptimal.
@@
Offline Trajectories vs. Simulated Rollouts. While both OPRL and Sim-OPRL depend on the offline dataset for estimating environment dynamics, they induce different suboptimality in modeling preference feedback. Simulated rollouts are designed to achieve good coverage of the optimal offline policy ${\pi}^{*}_{\textrm{offline}}$, which avoids wasting preference budget on trajectories with low rewards or high transition uncertainty. In contrast, as shown in Zhan et al. [2023a], due to the dependence of preferences on full trajectories, the reward concentrability term $C_{R}$ in Theorem 5.1 can be very large. While sampling from the offline buffer is not sensitive to the quality of the transition model, good coverage of the optimal policy is needed from both transition and preference data to achieve low suboptimality.
@@
Transition vs. Preference Model Quality. Our theoretical analysis also suggests an interesting trade-off in the sample efficiency of our approach, depending on the accuracy of the transition model. The width of the confidence interval reduces as significance parameter $\delta$ or dataset size increase, or as function class complexity $\mathcal{N}_{\mathcal{F}_{T}}$ decreases. For a target suboptimality gap $\epsilon$, provided the optimal offline policy $\pi^{*}_{\textrm{offline}}$ has a gap $\epsilon_{T}<\epsilon$, then the number of preferences required is of the order of $\mathcal{O}({\log(1/\delta)}/{(\epsilon-\epsilon_{T})^{2}})$. A more accurate transition model should therefore require fewer preference samples to achieve a given suboptimality, which we again confirm empirically.
6.3 Practical Implementation
We now complete the general algorithmic framework discussed above with a possible implementation strategy, allowing for empirical validation. In fact, with minor changes to the following framework, our paper also proposes a feasible implementation of related theoretical algorithms [Chen et al., 2022, Zhan et al., 2023a]. We refer the reader to Appendix C for further detail.
Model Learning and Policy Optimization.
Following prior work in offline reinforcement learning [Yu et al., 2020], we train ensembles of $N_{T}$ and $N_{R}$ neural network models for the transition and reward functions on different bootstraps of the data [Lakshminarayanan et al., 2017], denoted $\{\hat{T}_{1},\ldots\hat{T}_{N_{T}}\}$ and $\{\hat{R}_{1},\ldots\hat{R}_{N_{R}}\}$. We estimate MLE and uncertainty functions as follows:
$\hat{T}(\cdot|s,a)=\frac{1}{N_{T}}\sum_{i=1}^{N_{T}}\hat{T}_{i}(\cdot|s,a);$ $\quad u_{T}(s,a)=\max_{i,j\in[\![1,N_{T}]\!]}|\hat{T}_{i}(\cdot|s,a)-\hat{T}_{j}(\cdot|s,a)|_{1}\cdot R_{\textrm{max}}$
$\hat{R}(s,a)=\frac{1}{N_{R}}\sum_{i=1}^{N_{T}}\hat{R}_{i}(s,a);$ $\quad u_{R}(s,a)=\max_{i,j\in[\![1,N_{R}]\!]}|\hat{R}_{i}(s,a)-\hat{R}_{j}(s,a)|_{1}$
Each $\hat{R}_{i}$ in the ensemble has an associated preference function defined by the Bradley-Terry model, with $\sigma$ as the sigmoid function. We obtain preference uncertainty through variation over the ensemble as in Equation 2. Recall that transition and reward models are trained on $\mathcal{D}_{\textrm{offline}}$ and $\mathcal{D}_{\textrm{pref}}$ respectively; for computational efficiency, we sample preferences in batches of to reduce the number of reward model updates needed.
We approximate the pessimistic objective in Equation 3 by penalizing the reward function with the uncertainty, as in Lagrangian formulations of model-based offline RL [Yu et al., 2020, Rigter et al., 2022]. We solve for the following objective with a traditional reinforcement learning algorithm:
$\hat{\pi}^{*}=\text{argmax}_{\pi\in\Pi}\mathbb{E}_{(s,a)\sim d^{\pi}_{\hat{T}}}[\hat{R}(s,a)-\lambda_{R}u_{R}(s,a)-\lambda_{T}u_{T}(s,a)],$ (5)
where hyperparameters $\lambda_{T},\lambda_{R}$ control the degree of conservatism. Note that in our theoretical analysis, this was achieved through parameters $\beta_{T},\beta_{R}$ which affect the width of the confidence intervals $u_{R}$ and $u_{T}$, but their exact value cannot be estimated. We show in Section A.2 that Equation 5 indeed lower bounds the true value function, under well-calibrated uncertainty estimates.
Near-Optimal Policy Set and Exploratory Policies.
Sim-OPRL requires constructing $\Pi_{\textrm{offline}}$, a set of near-optimal policies within a pessimistic model of the environment. Following Lindner et al. [2021], we obtain a policy model for each element $\hat{R}_{i}$ of the reward ensemble. Policy models are optimized to maximize returns under the transition model $\hat{T}$ and the reward function $\hat{R}_{i}-\lambda_{T}u_{T}$, ensuring pessimism w.r.t transitions. Next, the most exploratory policies are identified by generating rollouts of each candidate policy within the learned model $\hat{T}$. The trajectories $(\tau_{1},\tau_{2})$ induced by different policies and maximizing the preference uncertainty function $u_{P_{R}}(\tau_{1},\tau_{2})$ are used for preference feedback. We refer the reader to Appendix C for further detail.
@@ 7 Experimental results
In this section, we demonstrate the effectiveness of our preference elicitation strategy, Sim-OPRL, across a range of offline reinforcement learning environments and datasets. We demonstrate its superior performance over OPRL, as expected from our theoretical analysis.
@@
Since our closest related works do not propose any experimental validation [Chen et al., 2022, Zhan et al., 2023a], we propose a practical implementation of Preference-based Optimistic Planning (PbOP) in Appendix C; this elicitation method queries feedback over trajectory rollouts in the true environment [Chen et al., 2022]. We also compare against OPRL [Shin et al., 2022] with uniform and uncertainty-sampling. Finally, we report the performance of $\pi^{*}_{\textrm{offline}}$ and $\pi^{*}$ as upper bounds for the performance of our algorithm: the former is trained in the learned transition model with access to the true reward, and the latter has full knowledge of both transition and reward function.
We compare different preference elicitation strategies on a range of environments detailed in Appendix D. Among others, we explore environments from the D4RL benchmark [Fu et al., 2020] identified as particularly challenging offline preference-based reinforcement learning tasks [Shin et al., 2022], as well as a medical simulation designed to model the evolution of patients with sepsis [Oberst and Sontag, 2019]. As detailed in Appendix D, these environments consist of high-dimensional state spaces with continuous or discrete action spaces, follow complex transition dynamics, and have sparse or non-linear rewards and termination conditions. This makes them representative of the challenge of learning a reward function and learning offline in a real-world application. In particular, the sepsis simulation environment is commonly used in medically-motivated offline RL work [Tang and Wiens, 2021, Pace et al., 2023], and highlights another advantage of Sim-OPRL over OPRL: it does not require feedback on real trajectories from the observational dataset. In a sensitive setting such as healthcare where data access is carefully controlled, it may be attractive to query experts about synthetic trajectories rather than real samples.
Performance against State-of-the-Art. Performance and sample complexity results with different preference elicitation methods are given in Figure 1 and Table 2. Within the offline approaches, Sim-OPRL consistently achieves better environment returns than OPRL with much fewer preference queries. In line with our theoretical analysis, our empirical results therefore demonstrate that policy-based sampling in Sim-OPRL is more efficient than maximizing information gain on the reward function (uncertainty-based OPRL), which echoes similar conclusions reached in prior work on online preference elicitation [Lindner et al., 2021, Chen et al., 2022].
@@
As an upper bound for the performance of our algorithm, we include baselines that have access to the environment: we report the performance of the optimal policy $\pi^{*}$, as well as that of an algorithm querying feedback over optimistic rollouts in the real environment [Chen et al., 2022, PbOP]. In Figure 1, the PbOP method naturally reaches a superior policy with fewer samples as it allows environment interaction and can thus improve its estimate of the transition model in parallel to learning the preference function. As supported by our theoretical analysis, this result stresses the importance of having a high-quality transition model to make our method effective. We explore this in more detail in our following ablations.
Table 2: Sample complexity $N_{p}$ under different preference elicitation strategies, to reach a suboptimality gap of $\epsilon=20$ over normalized returns. Mean and 95% confidence interval over 6 experiments. The best-performing offline method is highlighted in bold,    marks when the target suboptimality could not be achieved. Note that PbOP has an advantage by having access to direct interaction with the environment.
Star MDP 32 $\pm$ 4 30 $\pm$ 4 4 $\pm$ 2 4 $\pm$ 2
Gridworld 105 $\pm$ 11 66 $\pm$ 7 49 $\pm$ 7 32 $\pm$ 4
MiniGrid-FourRooms 92 $\pm$ 7 53 $\pm$ 5 41 $\pm$ 5 25 $\pm$ 3
HalfCheetah-Random 108 $\pm$ 9 71 $\pm$ 8 50 $\pm$ 10 36 $\pm$ 3
Sepsis Simulation 642 $\pm$ 72 225 $\pm$ 46 75 $\pm$ 11
(a) StarMDP.
(b) Gridworld.
(c) MiniGrid-FourRooms.
(d) HalfCheetah-Random.
(e) Sepsis Simulation.
@@
Figure 1: Environment returns under different preference elicitation strategies. Mean and 95% confidence interval over 6 experiments. Environment returns are normalized between 0 and 100. Only OPRL and Sim-OPRL are fully offline.
Figure 2: Algorithm ablations (StarMDP).
Algorithm Ablations. We conduct ablations for our algorithm on a simple tabular MDP, with results in Figure 2. This example (transition and reward details deferred to Appendix D) illustrates the importance of pessimism with respect to the transition model. Even with access to true rewards, $\pi^{*}_{\textrm{offline}}$ is pessimistic to avoid the out-of-distribution state, as it is unclear how to reach it. Thus, in Figure 2, we see a drop in performance if pessimism is not applied to the output policy (purple lines). This confirms the theoretical insights from Zhu et al. [2023], Zhan et al. [2023a], who demonstrate the importance of pessimism in offline preference-based RL problems. Pessimism is also crucial in simulated rollouts, to avoid wasting preference budget on regions of low confidence — as value estimates are inaccurate in any case. This is reflected in lower performance without pessimism w.r.t the transition model in Figure 2 (brown line), and which could be seen as the naive adaptation of online preference elicitation methods to our setting [Chen et al., 2022, Lindner et al., 2022]. We also note the importance of optimism with respect to the reward uncertainty, both in OPRL in Figure 1 and in our own model-based rollouts in Figure 2.
@@
Transition vs. Preference Model Quality. Next, we empirically study the trade-off between transition and preference model performance in our problem setting. Still in the Star MDP, in the low-data regime, the error $\epsilon_{T}$ incurred in estimating the value function due to the misspecification of the transition model is large. As dictated by our theoretical analysis and as visualized in Figure 3(a), this significantly increases the number of preference samples $N_{p}$ required to achieve good final performance. At the other end of the spectrum, if the offline dataset is large and allows modeling the transition model accurately, then $\epsilon_{T}$ is small and the number of preference samples $N_{p}$ needed shrinks. We observe a similar trend for both Sim-OPRL and our OPRL uncertainty-sampling baseline.